OpenStudy (anonymous):

Consider the function \(f(x) = \tan(x) := \tan(x) := \frac{sin(x)}{cos(x)}\) a) Determine the maximal Interval of deinfintion \(I\) and \(0\)

OpenStudy (zarkon):

you want the maximum interval ,that contains zero, on which f(x) can be defined?

OpenStudy (anonymous):

yes..

OpenStudy (zarkon):

find where \(\cos(x)=0\)

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

bu how..

OpenStudy (zarkon):

\(\cos(x)\) and \(\sin(x)\) are continuous functions so the only place \(f(x)\) can't be defined is when the denominator is equal to zero

OpenStudy (zarkon):

refer to the unit circle

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

hm @Zarkon i need full solution, if i do it by myself i will get low points which could led me to be unable to participate at final exams :)

OpenStudy (anonymous):

and soon i must give up this homework, here its in germany 04:45 min and i must give it up at around 8 to 10 o clock

OpenStudy (anonymous):

ok thanks anyway :)

OpenStudy (zarkon):

sorry...I was taking a shower